OpenStudy (anonymous):

Solve using substitution: 4x − y − 4 = 0 7y = 29 + 9x

OpenStudy (anonymous):

Solve using substitution: -x + 2y = -7 10y − 5x = 21 Solve using substitution: 4x − y − 4 = 0 7y = 29 + 9x Solve using substitution: -y − 5x = -7 2y − 2x = 14 this about all i can do

OpenStudy (anonymous):

do you still need help or do you have it